OpenStudy (anonymous):

Determine the equations of both lines that are tangent to the graph of f(x)=x^2 and pass through point (1,-3).

OpenStudy (anonymous):

I got the slope: m=2

OpenStudy (anonymous):

(1, -3) is not a point on the graph..

OpenStudy (phi):

tangent to the graph of f(x)=x^2 means one point on the line is also on this curve, at (x, x^2) because y= x^2 the slope of the tangent line will be 2x (take the derivative of x^2) use the point slope formula through point (1,-3) y - (-3) = m ( x - 1) slope is 2x, y is x^2 x^2 +3 = 2x(x-1) simplify and solve for x (2 values) now you have two points: (1,-3) and (x,x^2) (evaluated at the correct x) from which you can find an equation of the tangent line

OpenStudy (anonymous):

aahh.. got it thanks a lot!!!(^_^)